find the values of m such that this question has no solution sets:
4x-2y=12
3x+my=2

Answer by longjonsilver(2297)

no solution means the 2 lines will not cross (the crossing point is the solution). This only occurs when the lines are parallel.

So, work out gradient of first line...

4x-2y=12
-2y=-4x+12
y = 2x - 6 --> gradient = 2

so second line needs "2x", so: 3x+my=2 gives

my = -3x+2
y = (-3/m)x + 2/m

so, -3/m has to be equal to 2...so

-3/m = 2
--> m = -3/2

jon.